OpenStudy (anonymous):

What is the solution to the open sentence? 4y – 3 = y + 6

OpenStudy (anonymous):

4y - 3 = y + 6 Add 3 to both sides of the equation: 4y = 1y + 9 Subtract 1y from both sides of the equation: 3y = 9 Divide: y = 3

OpenStudy (petiteme):

4y – 3 = y + 6 3y=9 y=9/3 y=3 :)